### v1

# Generalized Schalkwijk-Kailath Coding for Autoregressive Gaussian Channels

Abstract
We propose a Gaussian random coding scheme for AR($p$) Gaussian channels that generalizes the celebrated Schalkwijk-Kailath (SK) coding scheme. This constructive coding scheme, termed the SK(2) coding scheme, yields a closed-form characterization for the corresponding achievable rate. Among many others, this result shows that the celebrated SK coding scheme is not universally optimal, and therefore, disprove the conjecture proposed by Butman in [2].
1 Introduction and Summary
We consider the following additive colored Gaussian noise (ACGN) channel with feedback
$Y_{i}=X_{i}(M,Y_{1}^{i-1})+Z_{i},\penalty 10000\ \penalty 10000\ i=1,2,$... (1)
where $M$ denotes the transmitted message, and $\{Z_{i}\}_{i=1}^{\infty}$ is a zero-mean stationary Gaussian noise process, independent of $M$. The channel input $X_{i}$ at time $i$ may causally depend on the message $M$ and the previous channel outputs $Y_{1}^{i-1}\triangleq\{Y_{1},...,Y_{i-1}\}$. And the channel input sequence $\{X_{i}\}_{i=1}^{\infty}$ is subject to an average power constraint $P>0$; i.e., for every $n$,
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[X^{2}_{i}(M,Y_{1}^{i-1})]\leq P.$
Let $C_{\mathrm{FB}}(P)$ denote the capacity of the ACGN channel (1), commonly termed the Gaussian feedback capacity in the literature.
It is well-known that the nonfeedback capacity of the channel (1), denoted by $C(P)$, can be be determined by the water-filling method [13]. In the special case where the noise process $\{Z_{i}\}$ is white, i.e., $Z_{i}=W_{i}$ with $\{W_{i}\}$ a white Gaussian process with zero mean and unit variance, the capacity of this additive white Gaussian noise (AWGN) channel was first determined by Shannon [15] (see also [18, 19]). Furthermore, Shannon [14] proved that for this channel, feedback does not increase capacity; that is, the feedback capacity $C_{\mathrm{FB}}(P)$ coincides with the nonfeedback capacity $C(P)$ (see also [7, 8]).
A significant later result, due to Schalkwijk and Kailath [12, 11], showed that a simple linear feedback coding scheme, commonly referred to as the SK coding scheme, achieves the capacity of the AWGN channel. For the general $p$th-order autoregressive (AR($p$)) Gaussian channel under average power constraint $P$, where the noise process $\{Z_{i}\}$ in (1) is a stationary AR(p)(p) Gaussian process, Butman [1] derived the optimal linear signaling structure with feedback and for stationary AR(1) Gaussian channels gave a closed-form lower bound on $C_{\mathrm{FB}}(P)$ (which coincides with $\bar{I}_{\mathrm{SK1}}(P)$ in Theorem 3.2 when $p=1$), achievable under equal signal energy allocation. Wolfowitz [17] provided an alternative derivation of the same bound and proved its optimality within a class of linear signaling schemes. This lower bound is in fact tight, as shown in [9, 21]. Upper bounds on $C_{\mathrm{FB}}(P)$ of the AR($p$) Gaussian channel were further provided in [16]. More recently, Kim in [9, Section 6] showed that the SK coding scheme achieves the feedback capacity for the first-order autoregressive moving-average (ARMA(1,1)) Gaussian channel. Based on a state-space approach, Kim further stated that a $k$-dimensional variant of the SK coding scheme achieves the feedback capacity for any ARMA noise spectrum of order kk. [^1] As noted in a recent paper [3], there is a gap in the proof of a key intermediate result [9, Corollary 4.4]; consequently, the proof of Theorem 6.1 in [9] is incomplete, leaving the optimality of the scheme as a conjecture.
In this paper, we consider the stationary AR($p$) Gaussian channel
$Y_{i}=X_{i}+Z_{i},\penalty 10000\ \penalty 10000\ i\geq 1,$ (2)
where the noise process $\{Z_{i}\}$ is a real-valued AR($p$) Gaussian process with parameters $\beta_{k}$, $|\beta_{k}|<1$ for all k=1,2,,pk=1,2,...,p, namely, it has the power spectral density [10]
$S_{Z}(e^{i\theta})=|H_{Z}(e^{i\theta})|^{2}=\frac{1}{|L_{Z}(e^{i\theta})|^{2}}=\frac{1}{|\prod_{k=1}^{p}(1+\beta_{k}e^{i\theta})|^{2}}.$ (3)
As mentioned above, the literature on optimal coding for stationary ARMA Gaussian channels is extensive. However, to the best of our knowledge, determining the optimal coding scheme and the corresponding feedback capacity, either analytically or numerically, remains an open problem for all ARMA orders beyond the ARMA(1,1) case. The main contribution of this work is a closed-form expression for the maximum achievable rate of a Gaussian random coding scheme, termed the SK(2) coding scheme, over the AR($p$) channel (2). This rate, denoted by $\bar{I}_{\mathrm{SK2}}(P)$, provides a lower bound on $C_{\mathrm{FB}}(P)$. We further show that the SK(2) coding scheme can strictly outperform the celebrated SK coding scheme, thereby disproving the conjecture of Butman [2].
The remainder of the paper is organized as follows. In Section 2, we introduce necessary notation and terminologies. Section 3 reviews the classical SK coding scheme and introduces the SK(1) coding scheme with its maximal achievable rate given in Theorem 3.2. Section 4 contains the main contributions:
  • $\bullet$
    We propose the SK(2) coding scheme for AR($p$) Gaussian channels, some type of generalized SK coding, and derive its achievable rate (Theorem 4.1, Corollary 4.2).
  • $\bullet$
    For AWGN and AR(1) Gaussian channels, we show that the SK(2) coding scheme achieves feedback capacity (Corollary 4.3 and Corollary 4.5).
  • $\bullet$
    For AR(2) Gaussian channels, the SK(2) coding scheme can strictly outperform the celebrated SK coding scheme, thereby disproving a conjecture of Butman [2] (Example 4.6).
2 Notation and Terminologies
We use $(\Omega,\mathcal{F},\mathbb{P})$ to denote the underlying probability space, and $\mathbb{E}$ to denote the expectation with respect to the probability measure $\mathbb{P}$. Let $X,Y$ be random variables defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$, which will be used to illustrate most of the notions and facts in this section (the same notations may have different connotations in other sections). In this work, a random variable may be real-valued (with a probability density function) or vector-valued (more precisely, taking values in $\mathbb{R}^{\infty}$ or $\mathbb{R}^{n}$). We use $\mu_{X}$ and $\mu_{Y}$ to denote the induced probability distributions of $X$ and $Y$, respectively. Their product distribution is denoted by $\mu_{X}\times\mu_{Y}$, and their joint distribution is $\mu_{XY}$. Throughout the remainder of this paper, unless otherwise specified, $\{W_{i}\}$ denotes a white Gaussian process with zero mean and unit variance.
The mutual information $I(X;Y)$ between two random variables $X,Y$ is defined as
$I(X;Y)=\begin{cases}{\mathbb{E}}\left[\log\frac{d\mu_{XY}}{d\mu_{X}\times\mu_{Y}}(X,Y)\right],&\mbox{ if }\frac{d\mu_{X{Y}}}{d\mu_{X}\times\mu_{Y}}\mbox{ exists},\\ \infty,&\mbox{otherwise},\end{cases}$
where $d\mu_{X{Y}}/d\mu_{X}\times\mu_{Y}$ denotes the Radon-Nikodym derivative of $\mu_{XY}$ with respect to $\mu_{X}\times\mu_{Y}$.
We use $\delta_{ij}$ to denote the Kronecker delta function. For $a\in\mathbb{C}$, $|a|$ denotes its modulus; for $\in\mathbb{C}^{n}, a\|a\|$ denotes its $\ell^{2}$-norm. For an $m\times n$ matrix $A$, we use $A^{\ast}$ to denote its conjugate transpose (or $A^{T}$ if $A$ is real). If $A$ is square, $|A|$ denotes its determinant, with the convention that $|A|=1$ for an empty matrix (i.e., $m=n=0$).
3 The Schalkwijk–Kailath Coding
It is well-known that the SK coding scheme achieves capacity for the AWGN channel with noiseless feedback under an average power constraint $P$. In the spirit of the exposition in [5, Section 17.1.1], the classical SK coding scheme can be reformulated in the following. Consider the discrete-time AWGN channel
$Y_{i}=X_{i}+W_{i},\penalty 10000\ \penalty 10000\ i\geq 1,$ (4)
where $\{W_{i}\}$ a white Gaussian process with zero mean and unit variance and $\{X_{i}\}$ are subject to the average power constraint $P$, i.e.,
$\sum_{i=0}^{n-1}\mathbb{E}[X_{i}^{2}]\leq nP.$
The encoding scheme proceeds as follows. At time $i=0$, the encoder transmits a symbol $X_{0}$ chosen arbitrarily from the interval $\sqrt{P},\sqrt{P}]$, yielding the output $Y_{0}=X_{0}+W_{0}$. Because of noiseless feedback, the encoder learns $Y_{0}$ and can compute the noise realization $W_{0}=Y_{0}-X_{0}$. At time $i=1$, the encoder transmits $X_{1}=\gamma_{1}W_{0}$, where $\gamma_{1}=\sqrt{P}$ ensures that $\mathbb{E}[X_{1}^{2}]=P$. For each subsequent time $i\geq 2$, the encoder computes the minimum mean-square error (MMSE) estimate $\mathbb{E}[W_{0}|Y_{1}^{i-1}]$ based on all previous channel outputs $Y_{1}^{i-1}$ available via feedback. It then transmits
$X_{i}=\gamma_{i}(W_{0}-\mathbb{E}[W_{0}|Y_{1}^{i-1}]),$ (5)
where $\gamma_{i}$ is chosen such that $\mathbb{E}[X_{i}^{2}]=P$ for all $i$. The SK coding scheme, given in (5), admits a recursive interpretation in which the transmitter repeatedly refines the receiver’s knowledge of the initial noise $W_{0}$, equivalently, the original message $X_{0}$, across successive uses of the channel. Specifically, the encoding rule can be written as
$X_{i}$ $=\gamma_{i}(W_{0}-\mathbb{E}[W_{0}|Y_{1}^{i-1}])$
$=\frac{\gamma_{i}}{\gamma_{i-1}}\left(X_{i-1}+\mathbb{E}[X_{i-1}|Y_{1}^{i-1}]\right)$
so that each transmission corrects the receiver’s estimation error incurred in the previous transmission.
Furthermore, define the process $V_{i}\triangleq\gamma_{i}W_{0}$. Since $W_{0}$ is independent of the forward noise $\{W_{i}\}_{i\geq 1}$, $\{V_{i}\}$ may be regarded as the underlying message process. Under this representation, the SK rule takes the following form
$X_{i}=V_{i}-\mathbb{E}[V_{i}|Y_{1}^{i-1}],$ (6)
where $\{V_{i}\}$ follows the deterministic recursion
$V_{i}=\frac{\gamma_{i}}{\gamma_{i-1}}V_{i-1},\penalty 10000\ \penalty 10000\ i\geq 2$ (7)
with initial condition $V_{1}=\sqrt{P}W_{0}$. It has been shown [4, 5, 11, 12] that for any such sequence $\{\gamma_{i}\}_{i\geq 1}$, this SK coding scheme (5), or equivalently, the recursive form (6) and (7), achieves the feedback capacity of the AWGN channel.
We now introduce a specialized instance of the SK coding scheme, which we term the SK(1) coding scheme. The name reflects that the underlying message process follows a first-order deterministic recursion. It is obtained from the recursive form in (6) and (7) by imposing the simple ratio condition
$\frac{\gamma_{i}}{\gamma_{i-1}}=\gamma\penalty 10000\ \penalty 10000\ \mbox{for all}\penalty 10000\ i,$
where $|\gamma|>1$. Under this condition, the message process $V=\{V_{i}\}$ follows a first-order deterministic recursion. Formally, the SK(1) coding scheme is defined by
$X_{i}=V_{i}-\mathbb{E}[V_{i}|Y_{1}^{i-1}]$ (8)
with the message process $\{V_{i}\}$ such that
$\begin{cases}V_{i+1}=\gamma V_{i},&i\geq 1,\\ V_{1}=U,\end{cases}$ (9)
where $U$ is a standard Gaussian random variable, independent of the noise process $Z=\{Z_{i}\}$. We define the corresponding SK(1) channel capacity as
$C_{\mathrm{SK1}}(P)=\sup_{(V,X)}\bar{I}(V;Y),$ (10)
where $\bar{I}(V;Y)$ denotes the mutual information rate between the processes $\{V_{i}\}$ and $\{Y_{i}\}$, and the supremum is taken over all pairs $(V,X)$ satisfying (8) and (9) under the asymptotic average power constraint
[…]
Note that the SK(1) coding scheme, which can be directly transformed into the original constructive SK coding, has been established as capacity-achieving for stationary ARMA(1,1) Gaussian channels with feedback [9].
We begin by stating a key lemma, the proof of which is deferred to Appendix A. This lemma characterizes the quadratic form associated with the inverse of a structured matrix.
Lemma 3.1.
Let $\{u_{k};k=1,2,...,m\}\subset\mathbb{C}^{n}$ for $n,m\geq 1$. Then, it holds that
$u_{i}^{\ast}(I_{n}+UU^{\ast})^{-1}u_{j}=\left\{\begin{aligned} 1\penalty 10000\ -\penalty 10000\ \frac{M_{ii}}{|A|}&\qquad\mbox{if $i=j$};\\ (-1)^{i+j+1}\frac{M_{ji}}{|A|}&\qquad\mbox{if $i\neq j$},\end{aligned}\right$.
where $U=\begin{bmatrix}u_{1}&u_{2}&...&u_{m}\end{bmatrix}$ and $A=I_{m}+U^{\ast}U$ and $M_{ij}$ is the $(i,j)$-minor of $A$, i.e., the determinant of the $(m-1)\times(m-1)$ matrix that results from deleting row $i$ and column $j$ of $A$. In particular, if $m=1$, then we have
$u_{1}^{\ast}(I_{n}+UU^{\ast})^{-1}u_{1}=1-1/A.$
The following theorem provides the maximal achievable rate $\bar{I}_{\mathrm{SK1}}(P)$ for the SK(1) coding scheme over stationary AR(p)(p) Gaussian channels, yielding a lower bound on $C_{\mathrm{SK1}}(P)$.
Theorem 3.2.
Suppose the noise $\{Z_{i}\}$ is an $AR(p)$ Gaussian process with parameters $\beta_{k}$, $|\beta_{k}|<1$ for all k=1,2,,pk=1,2,...,p, namely, it has the power spectral density
$S_{Z}(e^{i\theta})=|H_{Z}(e^{i\theta})|^{2}=\frac{1}{|L_{Z}(e^{i\theta})|^{2}}=\frac{1}{|\prod_{k=1}^{p}(1+\beta_{k}e^{i\theta})|^{2}}.$
Then, the maximal achievable rate for the SK(1) coding scheme $\bar{I}_{\mathrm{SK1}}(P)$ is given by
$\bar{I}_{\mathrm{SK1}}(P)\triangleq\max_{\gamma}\log|\gamma|,$ (12)
where the maximum is taken over all $|\gamma|>1$ satisfying the power constraint
$(\gamma^{2}-1)H^{2}_{Z}(\gamma^{-1})=P.$ (13)
Remark 3.3.
For the special cases $p=0$ and $p=1$, corresponding to the AWGN channel and AR(1) Gaussian channel, respectively, the feedback capacity $C_{\mathrm{FB}}(P)$ under power constraint $P$ is already known from existing literature (see, e.g., [9, 21]). It follows directly from these results that $C_{\mathrm{FB}}(P)=\bar{I}_{\mathrm{SK1}}(P)$, which implies that $C_{\mathrm{SK1}}(P)=C_{\mathrm{FB}}(P)$. This equivalence establishes the optimality of the SK(1) coding scheme the AWGN channel and AR(1) Gaussian channel.
Proof.
The proof follows a similar approach to that of Theorem 4.1 and is therefore deferred to Appendix B. ∎
To conclude this section, we have introduced the $\mathrm{SK}(1)$ coding scheme and derived an explicit expression for the maximal rate it achieves, as presented in Theorem 3.2. This result extends the formula for the SK(1) coding scheme from AR(1) to AR($p$) Gaussian channels. In the following section, we introduce the SK(2) coding scheme, wherein the underlying message process follows a second-order deterministic recursion, thereby generalizing the $\mathrm{SK}(1)$ coding scheme.
4 The SK(2) Coding Scheme
Building upon the framework of the SK(1) coding scheme, we now present its natural extension: the SK(2) coding scheme. This generalization is achieved by generalizing the underlying message process from a first-order to a second-order linear deterministic recursion. Formally, let $U_{1}$and $U_{2}$ be two independent standard normal random variables, independent of the noise process $\{Z_{i}\}$. The transmitter initially sends
$X_{1}=U_{1},\penalty 10000\ X_{2}=U_{2}$
and subsequently sends
$X_{i}=V_{i}-\mathbb{E}{[V_{i}|Y_{1}^{i-1}]},\penalty 10000\ i=3,4,$..., (14)
where $\{V_{n}\}$ is the message process satisfying
$\left\{\begin{aligned} V_{i+1}&=\penalty 10000\ aV_{i}+bV_{i-1},\penalty 10000\ \penalty 10000\ i\geq 2,\\ V_{2}\penalty 10000\ \penalty 10000\ &=\penalty 10000\ U_{2},\\ V_{1}\penalty 10000\ \penalty 10000\ &=\penalty 10000\ U_{1},\end{aligned}\right$. (15)
where $a,b\in\mathbb{R}$. Analogous to the definition of the SK(1) channel capacity in (10), we define the SK(2) channel capacity as
$C_{\mathrm{SK2}}(P)=\sup_{(V,X)}\bar{I}(V;Y),$
where the supremum is taken over all pairs $(V,X)$ satisfying (14) and (15) subject to the constraint (11).
It is clear to see that when $b=0$, the recursion (15) reduces to $V_{n+1}=aV_{n}$, essentially recovering the SK(1) coding scheme (9). Thus, the case of interest is $b\neq 0$. As with the SK(1) coding scheme, parameters $a$ and $b$ must satisfy certain constraints. Rather than imposing conditions directly, we analyze the corresponding characteristic equation (in $\lambda$) of (15):
$\lambda^{2}-a\lambda-b=0.$ (16)
Let $\gamma_{1},\penalty 10000\ \gamma_{2}$ denote its roots. Two cases arise:
  1. 1.
    Case A: $\gamma_{1}\neq\gamma_{2}$ (distinct roots);
  2. 2.
    Case B: $\gamma_{1}=\gamma_{2}=:\gamma$ (repeated root).
In either case, $\{V_{n}\}$ admits the finite-rank representation
$V_{n}=a_{n}U_{1}+b_{n}U_{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ n\geq 1,$
where the deterministic coefficients $(a_{n},b_{n})$ are given by:
  • $\bullet$
    Case A:
    $a_{n}=-(c_{1}\gamma_{2}\gamma_{1}^{n}+c_{2}\gamma_{1}\gamma_{2}^{n}),\penalty 10000\ b_{n}=c_{1}\gamma_{1}^{n}+c_{2}\gamma_{2}^{n},$ (17)
    with $c_{1},c_{2}$ defined as
    $c_{1}=\frac{1}{\gamma_{1}(\gamma_{1}-\gamma_{2})},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ c_{2}=\frac{1}{\gamma_{2}(\gamma_{2}-\gamma_{1})}.$
  • $\bullet$
    Case B:
    $a_{n}=(2-n)\gamma^{n-1},\penalty 10000\ b_{n}=(n-1)\gamma^{n-2}.$ (18)
Clearly, $a$ and $b$ are uniquely determined by $\gamma_{1}$ and $\gamma_{2}$; hence, it suffices to constrain the pair $(\gamma_{1},\gamma_{2})$. As shown in the following theorem, it is natural to require that $|\gamma_{1}|\wedge|\gamma_{2}|>1$.
We now state the corresponding achievable rate $\bar{I}_{\mathrm{SK2}}(P)$ for the SK(2) coding scheme under the condition $b\neq 0$ in (15) over stationary AR(p)(p) Gaussian channels, yielding a lower bound on $C_{\mathrm{SK2}}(P)$.
Theorem 4.1.
Suppose the noise $\{Z_{i}\}$ is an $AR(p)$ Gaussian process with parameters $\beta_{k}$, $|\beta_{k}|<1$ for all k=1,2,,pk=1,2,...,p, namely, it has the power spectral density
$S_{Z}(e^{i\theta})=|H_{Z}(e^{i\theta})|^{2}=\frac{1}{|L_{Z}(e^{i\theta})|^{2}}=\frac{1}{|\prod_{k=1}^{p}(1+\beta_{k}e^{i\theta})|^{2}}.$
Then, under the condition $b\neq 0$ in (15), the maximal achievable rate for the SK(2) coding $\bar{I}_{\mathrm{SK2}}(P)$ is given by
$\bar{I}_{\mathrm{SK2}}(P)\triangleq\max_{\gamma_{1},\gamma_{2}}2\log|\gamma_{1}|\wedge|\gamma_{2}|$ (19)
where the maximum is taken over all $|\gamma_{1}|>1,\penalty 10000\ |\gamma_{2}|>1$ satisfying the power constraint
$\frac{1}{\Delta}\left(\frac{L_{11}}{\gamma^{2}_{2}}+\frac{L_{22}}{\gamma_{1}^{2}}+\frac{2L_{12}}{\gamma_{1}\gamma_{2}}\right)\leq P,$ (20)
where
$\Delta$ $=(\gamma_{1}-\gamma_{2})^{2}(L_{11}L_{22}-L_{12}^{2}),$ (21)
$L_{ij}$ $=\frac{c_{i}c_{j}L_{Z}(\gamma_{i}^{-1})L_{Z}(\gamma_{j}^{-1})}{\gamma_{i}\gamma_{j}-1}\penalty 10000\ \penalty 10000\ \mbox{for $i,j=1,2,$}$
$c_{1}$ $=\frac{1}{\gamma_{1}(\gamma_{1}-\gamma_{2})}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ c_{2}=\frac{1}{\gamma_{2}(\gamma_{2}-\gamma_{1})}.$
Proof.
Consider the SK(2) coding scheme formulated by (14) and (15) over the stationary AR(p)(p) Gaussian channel $Y_{i}=X_{i}+Z_{i}$ with the noise spectral density function $S_{Z}(e^{i\theta})$ given by (3).
We first address Case A and the proof for Case B is deferred to the end of this section. Recall that
$Y_{n}$ $=X_{n}+Z_{n}$
$=V_{n}-\mathbb{E}[V_{n}|Y_{1}^{n-1}]+Z_{n}$
and define
$Y_{n}^{\ast}=V_{n}+Z_{n}.$
Clearly, $Y_{n}^{\ast}=Y_{n}+\mathbb{E}[V_{n}|Y_{1}^{n-1}]$ and therefore, for any $i\geq 1$, $\{Y_{1}^{i}\}$ and $\{Y_{1}^{\ast,i}\}$ are uniquely determined by each other. Furthermore, define
$\left\{\begin{aligned} \widetilde{Y}_{n}&=Y_{n}^{\ast},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ n\leq p\\ \widetilde{Y}_{n}&=L(B)Y^{\ast}_{n},\penalty 10000\ \penalty 10000\ \penalty 10000\ n\geq p+1\end{aligned}\right$. (22)
where $L$ stands for $L_{Z}$, $B$ denotes the backward‑shift operator, and both notations are used throughout the rest of the proof. Now by (17) and (18), we can express
$\widetilde{Y}_{n}$ $=L(B)(V_{n}+Z_{n})$ (23)
$=L(B)(a_{n}U_{1}+b_{n}U_{2})+L(B)Z_{n}$
$=d_{1,n}U_{1}+d_{2,n}U_{2}+W_{n}$
for $n\geq p+1$, where
$\left\{\begin{aligned} d_{1,n}&=-c_{1}\gamma_{2}\gamma_{1}^{n}L_{Z}(\gamma_{1}^{-1})-c_{2}\gamma_{1}\gamma_{2}^{n}L_{Z}(\gamma_{2}^{-1})\\ d_{2,n}&=c_{1}\gamma_{1}^{n}L_{Z}(\gamma_{1}^{-1})+c_{2}\gamma_{2}^{n}L_{Z}(\gamma_{2}^{-1}).\end{aligned}\right$. (24)
The mutual information between $V_{1}^{n}$ and $Y_{1}^{n}$ for $n\geq p+1$ is given by
$\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})$ $=\frac{1}{n}I(U_{1},U_{2};Y_{1}^{n})$ (25)
$=\frac{1}{n}I(U_{1};Y_{1}^{n})+\frac{1}{n}I(U_{2};Y_{1}^{n}|U_{1})$
$\geq\frac{1}{n}I(U_{1};Y_{1}^{n})+\frac{1}{n}I(U_{2};Y_{1}^{n})$
$=\frac{1}{n}I(U_{1};\widetilde{Y}_{1}^{n})+\frac{1}{n}I(U_{2};\widetilde{Y}_{1}^{n})$
$\geq\frac{1}{n}I(U_{1};\widetilde{Y}_{p+1}^{n})+\frac{1}{n}I(U_{2};\widetilde{Y}_{p+1}^{n})$
$=-\frac{1}{2n}\log\mathbb{E}[|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{p+1}^{n}]|^{2}]$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -\frac{1}{2n}\log\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{p+1}^{n}]|^{2}].$
On the other hand, for $n\geq p+2$,
$\mathbb{E}[X_{n}^{2}]$ $=\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|Y_{1}^{n-1}]|^{2}]$ (26)
$=\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]$
$\leq E[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{p+1}^{n-1}]|^{2}].$
Without loss of generality, our analysis focuses only on the limiting behavior of the inequalities (25) and (26); therefore, it suffices to redefine the sequence $\{\widetilde{Y}_{k}\}$ in (22) by the form given in (23), taking the initial index $n=1$ rather than $n=p+1$:
$\widetilde{Y}_{n}=d_{1,n}U_{1}+d_{2,n}U_{2}+W_{n}$
for $n\geq 1$. Since $U_{1}$ and $\{\widetilde{Y}_{1}^{n}\}$ are jointly Gaussian, it holds that
$\mathbb{E}[|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{1}^{n}]|^{2}]=1-D_{1,n}^{T}(I_{n}+D_{1,n}D_{1,n}^{T}+D_{2,n}D_{2,n}^{T})^{-1}D_{1,n},$ (27)
where
$D_{1,n}=[d_{1,1}\penalty 10000\ d_{1,2}\penalty 10000\ ...\penalty 10000\ d_{1,n}]^{T},\penalty 10000\ D_{2,n}=[d_{2,1}\penalty 10000\ d_{2,2}\penalty 10000\ ...\penalty 10000\ d_{2,n}]^{T}.$
It then follows from Lemma 3.1 that (27) boils down to
$\mathbb{E}[|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{1}^{n}]|^{2}]=\frac{1+D_{2,n}^{T}D_{2,n}}{|I_{2}+D_{n}^{T}D_{n}|},$ (28)
where $I_{2}$ is the $2\times 2$ identity matrix and $D_{n}=[D_{1,n}\penalty 10000\ D_{2,n}]$. Similarly, we can establish
$\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{1}^{n}]|^{2}]=\frac{1+D_{1,n}^{T}D_{1,n}}{|I_{2}+D_{n}^{T}D_{n}|}.$ (29)
On the other hand, since $V_{n}$ and $\{\widetilde{Y}_{1}^{n-1}\}$ are also jointly Gaussian, it holds that
$\mathbb{E}$ $[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]=\text{Var}(V_{n})-\Sigma_{V_{n}|\widetilde{Y}_{1}^{n-1}}\Sigma_{\widetilde{Y}_{1}^{n-1}}^{-1}\Sigma_{\widetilde{Y}_{1}^{n-1}|V_{n}}$ (30)
$=a_{n}^{2}+b_{n}^{2}-\begin{bmatrix}a_{n}&b_{n}\end{bmatrix}D_{n-1}^{T}(I_{n-1}+D_{n-1}D_{n-1}^{T})^{-1}D_{n-1}\begin{bmatrix}a_{n}\\ b_{n}\end{bmatrix}$
$=\begin{bmatrix}a_{n}&b_{n}\end{bmatrix}(I_{2}-D_{n-1}^{T}(I_{n-1}+D_{n-1}D_{n-1}^{T})^{-1}D_{n-1})\begin{bmatrix}a_{n}\\ b_{n}\end{bmatrix}$
$=\begin{bmatrix}a_{n}&b_{n}\end{bmatrix}\begin{bmatrix}\frac{1+D_{2,n-1}^{T}D_{2,n-1}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}&-\frac{D_{1,n-1}^{T}D_{2,n-1}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}\\ -\frac{D_{2,n-1}^{T}D_{1,n-1}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}&\frac{1+D_{1,n-1}^{T}D_{1,n-1}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}\end{bmatrix}\begin{bmatrix}a_{n}\\ b_{n}\end{bmatrix}$
$=\frac{a_{n}^{2}+b_{n}^{2}+\|a_{n}D_{2,n-1}-b_{n}D_{1,n-1}\|^{2}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}.$
Now, we are ready to evaluate the rates of convergence of (28), (29) and (30) under Case A. Since $d_{1,n}$ and $d_{2,n}$ can be alternatively expressed as
$d_{1,n}$ $=\begin{bmatrix}-\gamma_{2}&-\gamma_{1}\end{bmatrix}\begin{bmatrix}c_{1}L(\gamma_{1}^{-1})\gamma_{1}^{n}\\[6.0pt] c_{2}L(\gamma_{2}^{-1})\gamma_{2}^{n}\end{bmatrix}$
$d_{2,n}$ $=\begin{bmatrix}1&1\end{bmatrix}\begin{bmatrix}c_{1}L(\gamma_{1}^{-1})\gamma_{1}^{n}\\[6.0pt] c_{2}L(\gamma_{2}^{-1})\gamma_{2}^{n}\end{bmatrix}$,
we have
$D_{1,n}^{T}D_{1,n}=\sum_{k=1}^{n}|d_{1,k}|^{2}$ (31)
$=\sum_{k=1}^{n}\begin{bmatrix}-\gamma_{2}&-\gamma_{1}\end{bmatrix}\begin{bmatrix}c_{1}L(\gamma_{1}^{-1})\gamma_{1}^{n}\\[6.0pt] c_{2}L(\gamma_{2}^{-1})\gamma_{2}^{n}\end{bmatrix}\begin{bmatrix}c_{1}L(\gamma_{1}^{-1})\gamma_{1}^{n}&c_{2}L(\gamma_{2}^{-1})\gamma_{2}^{n}\end{bmatrix}\begin{bmatrix}-\gamma_{2}\\[6.0pt] -\gamma_{1}\end{bmatrix}$
$=\sum_{k=1}^{n}\begin{bmatrix}\gamma_{2}&\gamma_{1}\end{bmatrix}\begin{bmatrix}c_{1}^{2}L^{2}(\gamma_{1}^{-1})\gamma_{1}^{2n}&c_{1}c_{2}L(\gamma_{1}^{-1})L(\gamma_{2}^{-1})\gamma_{1}^{n}\gamma_{2}^{n}\\[6.0pt] c_{1}c_{2}L(\gamma_{1}^{-1})L(\gamma_{2}^{-1})\gamma_{1}^{n}\gamma_{2}^{n}&c_{2}^{2}L^{2}(\gamma_{2}^{-1})\gamma_{2}^{2n}\end{bmatrix}\begin{bmatrix}\gamma_{2}\\[6.0pt] \gamma_{1}\end{bmatrix}$
$=\begin{bmatrix}\gamma_{2}&\gamma_{1}\end{bmatrix}\begin{bmatrix}L_{11}\gamma_{1}^{2}(\gamma_{1}^{2n}-1)&L_{12}\gamma_{1}\gamma_{2}(\gamma_{1}^{n}\gamma_{2}^{n}-1)\\[6.0pt] L_{12}\gamma_{1}\gamma_{2}(\gamma_{1}^{n}\gamma_{2}^{n}-1)&L_{22}\gamma_{2}^{2}(\gamma_{2}^{2n}-1)\end{bmatrix}\begin{bmatrix}\gamma_{2}\\[6.0pt] \gamma_{1}\end{bmatrix}$
$\triangleq\begin{bmatrix}\gamma_{2}&\gamma_{1}\end{bmatrix}L_{n}^{(2)}\begin{bmatrix}\gamma_{2}\\[6.0pt] \gamma_{1}\end{bmatrix}$,
where we have defined
$L_{n}^{(2)}=\begin{bmatrix}L_{11}\gamma_{1}^{2}(\gamma_{1}^{2n}-1)&L_{12}\gamma_{1}\gamma_{2}(\gamma_{1}^{n}\gamma_{2}^{n}-1)\\[6.0pt] L_{12}\gamma_{1}\gamma_{2}(\gamma_{1}^{n}\gamma_{2}^{n}-1)&L_{22}\gamma_{2}^{2}(\gamma_{2}^{2n}-1)\end{bmatrix}$.
Then, similarly, it holds that
$D_{2,n}^{T}D_{2,n}$ $=\begin{bmatrix}1&1\end{bmatrix}L_{n}^{(2)}\begin{bmatrix}1\\ 1\end{bmatrix}$ (32)
$D_{1,n}^{T}D_{2,n}$ $=\begin{bmatrix}-\gamma_{2}&\gamma_{1}\end{bmatrix}L_{n}^{(2)}\begin{bmatrix}1\\ 1\end{bmatrix}$,
whence,
$D_{1,n}^{T}D_{1,n}D_{2,n}^{T}D_{2,n}$- $D_{1,n}^{T}D_{2,n}D_{2,n}^{T}D_{1,n}=\sum_{i<j}(d_{1,i}d_{2,j}-d_{1,j}d_{2,i})^{2}$ (33)
$=(\gamma_{1}-\gamma_{2})^{2}c_{1}^{2}c_{2}^{2}\gamma_{1}^{2}\gamma_{2}^{2}L^{2}(\gamma_{1}^{-1})L^{2}(\gamma_{2}^{-1})\frac{(\gamma_{1}^{2n}-1)(\gamma_{2}^{2n}-1)}{(\gamma_{1}^{2}-1)(\gamma_{2}^{2}-1)}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -(\gamma_{1}-\gamma_{2})^{2}c_{1}^{2}c_{2}^{2}\gamma_{1}^{2}\gamma_{2}^{2}L^{2}(\gamma_{1}^{-1})L^{2}(\gamma_{2}^{-1})\frac{(\gamma_{1}^{n}\gamma_{2}^{n}-1)^{2}}{(\gamma_{1}\gamma_{2}-1)^{2}}$
$=\Delta\gamma_{1}^{2}\gamma_{2}^{2}(\gamma_{1}^{2n}\gamma_{2}^{2n}-1)+2L^{2}_{12}\gamma_{1}^{2}\gamma_{2}^{2}(\gamma_{1}-\gamma_{2})^{2}\gamma_{1}^{n}\gamma_{2}^{n}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -\frac{(\gamma_{1}\gamma_{2}-1)^{2}\Delta\gamma_{1}^{2}\gamma_{2}^{2}}{(\gamma_{1}-\gamma_{2})^{2}}(\gamma_{1}^{2n}+\gamma_{2}^{2n}).$
Thus,
$|I_{2}+D_{n}^{T}D_{n}|$ =|[$=\left|\begin{bmatrix}1+D_{1,n}^{T}D_{1,n}&D_{1,n}^{T}D_{2,n}\\ D_{2,n}^{T}D_{1,n}&1+D_{2,n}^{T}D_{2,n}\end{bmatrix}\right$| (34)
$=(1+D_{1,n}^{T}D_{1,n})(1+D_{2,n}^{T}D_{2,n})-D_{1,n}^{T}D_{2,n}D_{2,n}^{T}D_{1,n}$
$=1+D_{1,n}^{T}D_{1,n}+D_{2,n}^{T}D_{2,n}+D_{1,n}^{T}D_{1,n}D_{2,n}^{T}D_{2,n}-D_{1,n}^{T}D_{2,n}D_{2,n}^{T}D_{1,n},$
which, together with (31) (32) and (33), immediately implies the leading‑order approximation with coefficient
$|I_{2}+D_{n}^{T}D_{n}|\approx\Delta\gamma_{1}^{2}\gamma_{2}^{2}\gamma_{1}^{2n}\gamma_{2}^{2n}.$ (35)
As a result, consider the rates of convergence of (28) and (29). It turns out from (31) and (32) and (35) that
[…]
or equivalently,
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}\left[\left|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{1}^{n}]\right|^{2}\right]=\log(|\gamma_{1}|\wedge|\gamma_{2}|).$ (36)
Similarly, we also have
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}\left[\left|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{1}^{n}]\right|^{2}\right]=\log(|\gamma_{1}|\wedge|\gamma_{2}|).$ (37)
Therefore, combining (25), (36), and (37), it follows that the mutual information rate achieved by the SK(2) coding scheme specified in equations (14) and (15) can be lower bounded by
$\lim_{n\to\infty}\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})\geq 2\log(|\gamma_{1}|\wedge|\gamma_{2}|),$ (38)
which establishes the desired lower bound for the mutual information rate achieved by the SK(2) coding scheme.
Furthermore, we then check the limiting behavior of (30) as $n\to\infty$. To this end, we can easily see that
$a_{n}d_{2,k}-b_{n}d_{1,k}=\begin{bmatrix}\gamma_{2}^{n-1}&\gamma_{1}^{n-1}\end{bmatrix}\begin{bmatrix}c_{1}L(\gamma_{1}^{-1})\gamma_{1}^{k}\\[6.0pt] c_{2}L(\gamma_{2}^{-1})\gamma_{2}^{k}\end{bmatrix}$,
which implies that
$\|a_{n}D_{2,n-1}-b_{n}D_{1,n-1}\|^{2}=\begin{bmatrix}\gamma_{2}^{n-1}&\gamma_{1}^{n-1}\end{bmatrix}L_{n-1}^{(2)}\begin{bmatrix}\gamma_{2}^{n-1}\\[6.0pt] \gamma_{1}^{n-1}\end{bmatrix}$ (39)
$=L_{11}\gamma_{1}^{2}\gamma_{2}^{2(n-1)}(\gamma_{1}^{2(n-1)}-1)+L_{22}\gamma_{2}^{2}\gamma_{1}^{2(n-1)}(\gamma_{2}^{2(n-1)}-1)+2L_{12}\gamma_{1}\gamma_{2}\gamma_{1}^{n-1}\gamma_{2}^{n-1}(\gamma_{1}^{n-1}\gamma_{2}^{n-1}-1).$
Taking the limit as $n\to\infty$ under Case A, inequality (30) yields
$\varlimsup_{n\to\infty}\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}] lim$$\leq\varlimsup_{n\to\infty}\frac{a_{n}^{2}+b_{n}^{2}+\|a_{n}D_{2,n-1}-b_{n}D_{2,n-1}\|^{2}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}$
$=\frac{L_{11}\gamma_{1}^{2}+L_{22}\gamma_{2}^{2}+2L_{12}\gamma_{1}\gamma_{2}}{\Delta\gamma_{1}^{2}\gamma_{2}^{2}}$
$=\frac{1}{\Delta}\left(\frac{L_{11}}{\gamma_{2}^{2}}+\frac{L_{22}}{\gamma_{1}^{2}}+\frac{2L_{12}}{\gamma_{1}\gamma_{2}}\right)$
$\leq P,$
where the first inequality follows from (35) and (39). This establishes the power constraint (26) for the SK(2) coding scheme. Together with (38), this completes the proof for Case A.
Next, we proceed to establish the result for Case B. Recall that in this case we have $\gamma_{1}=\gamma_{2}=\gamma\in\mathbb{R}$ with $|\gamma|>1$, which yields $a_{n}=(2-n)\gamma^{n-1}$ and $b_{n}=(n-1)\gamma^{n-2}$ as in (18). In the same way we define $\{\widetilde{Y}\}$ as in (22) and thus for $n\geq p+1$
$\widetilde{Y}_{n}=d_{1,n}U_{1}+d_{2,n}U_{2}+W_{n},$
where
$d_{1,n}$ $=\left[(2-n)L(\gamma^{-1})+L^{\prime}(\gamma^{-1})\gamma^{-1}\right]\gamma^{n-1}$
$d_{2,n}$ $=\left[(n-1)L(\gamma^{-1})-L^{\prime}(\gamma^{-1})\gamma^{-1}\right]\gamma^{n-2}$
for $n\geq 1$. Consequently, the scalar products $D^{T}_{1,n}D_{1,n}$, $D^{T}_{2,n}D_{2,n}$ and $D^{T}_{1,n}D_{2,n}$ are readily computed as follows:
$D^{T}_{1,n}D_{1,n}$ $=L^{2}(\gamma^{-1})\gamma^{2}S^{(2)}_{n-1}-2L(\gamma^{-1})L^{\prime}(\gamma^{-1})\gamma S^{(1)}_{n-1}+L^{2}(\gamma^{-1})$ (40)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{(L^{\prime}(\gamma^{-1}))^{2}\gamma^{-2}}{\gamma^{2}-1}(\gamma^{2n}-1)+2L(\gamma^{-1})L^{\prime}(\gamma^{-1})\gamma^{-1},$
$D^{T}_{2,n}D_{2,n}$ $=L^{2}(\gamma^{-1})\gamma^{-2}S_{n}^{(2)}-2L(\gamma^{-1})L^{\prime}(\gamma^{-1})\gamma^{-3}S_{n}^{(1)}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{(L^{\prime}(\gamma^{-1}))^{2}\gamma^{-4}}{\gamma^{2}-1}(\gamma^{2n}-1),$
$D^{T}_{1,n}D_{2,n}$ $=L(\gamma^{-1})\left(L^{\prime}(\gamma^{-1})+L^{\prime}(\gamma^{-1})\gamma^{-2}-L(\gamma^{-1})\gamma\right)S_{n-1}^{(1)}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +L(\gamma^{-1})L^{\prime}(\gamma^{-1})\gamma^{-4}\gamma^{2n}(n-1)-L^{2}(\gamma^{-1})\gamma S^{(2)}_{n-1}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -\frac{(L^{\prime}(\gamma^{-1}))^{2}\gamma^{-3}}{\gamma^{2}-1}(\gamma^{2n}-1)-L(\gamma^{-1})L^{\prime}(\gamma^{-1})\gamma^{-2},$
where
$S^{(1)}_{n}$ $\triangleq\sum_{k=1}^{n-1}k\gamma^{2k}$ (41)
$=\frac{1}{\gamma^{2}-1}n\gamma^{2n}-\frac{\gamma^{4}}{(\gamma^{2}-1)^{2}}\gamma^{2(n-1)}+\frac{\gamma^{2}}{(\gamma^{2}-1)^{2}}$
and
$S^{(2)}_{n}$ $\triangleq\sum_{k=1}^{n-1}k^{2}\gamma^{2k}$ (42)
$=\frac{1}{\gamma^{2}-1}n^{2}\gamma^{2n}-\frac{2\gamma^{2}}{(\gamma^{2}-1)^{2}}n\gamma^{2n}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{\gamma^{4}(\gamma^{2}+1)}{(\gamma^{2}-1)^{3}}\gamma^{2(n-1)}-\frac{(\gamma^{2}+1)\gamma^{2}}{(\gamma^{2}-1)^{3}}.$
Proceeding analogously to (34) and (35), we derive
$|I_{2}+D_{n}^{T}D_{n}|$ $=(1+D^{T}_{1,n}D_{1,n})(1+D^{T}_{2,n}D_{2,n})-(D^{T}_{1,n}D_{2,n})^{2}$ (43)
$\approx\frac{L^{4}(\gamma^{-1})}{(\gamma^{2}-1)^{4}}\gamma^{4(n+1)}.$
Applying this together with (40) (41) and (42) to both (28) and (29), and subsequently taking the limit as $n\to\infty$ as in (36) and (37), leads to
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}[|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{1}^{n}]|^{2}]=\log|\gamma|$
and
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{1}^{n}]|^{2}]=\log|\gamma|,$
which immediately implies that, upon taking the limit $n\to\infty$, inequality (25) reduces to
[…]
This inequality provides the desired lower bound on the mutual information rate achieved by the SK(2) coding scheme in Case B.
On the other hand, we next proceed to evaluate the limit of the power usage as in (30). To see this, we first have
$a_{n}d_{2,k}-b_{n}d_{1,k}=\gamma^{n-3}(H^{(2)}_{n})^{T}\begin{bmatrix}k\\ 1\end{bmatrix}\gamma^{k},$
where
$H^{(2)}_{n}\triangleq\begin{bmatrix}L(\gamma^{-1})&-\left(nL(\gamma^{-1})+\gamma^{-1}L^{\prime}(\gamma^{-1})\right)\end{bmatrix}$.
This implies that
$\|a_{n}D_{2,n-1}-b_{n}D_{1,n-1}\|^{2}$ $=\sum_{k=1}^{n-1}(a_{n}d_{2,k}-b_{n}d_{1,k})^{2}$ (44)
$=\gamma^{2(n-3)}\sum_{k=1}^{n-1}(H^{(2)}_{n})^{T}\begin{bmatrix}k\\ 1\end{bmatrix}\begin{bmatrix}k&1\end{bmatrix}H^{(2)}_{n}\gamma^{2k}$
$=(H^{(2)}_{n})^{T}\begin{bmatrix}\sum_{k=1}^{n-1}k^{2}\gamma^{2k}&\sum_{k=1}^{n-1}k\gamma^{2k}\\[6.0pt] \sum_{k=1}^{n-1}k\gamma^{2k}&\sum_{k=1}^{n-1}\gamma^{2k}\end{bmatrix}H^{(2)}_{n}\gamma^{2(n-3)}$
$=\gamma^{2(n-3)}(H^{(2)}_{n})^{T}\begin{bmatrix}S_{n}^{(2)}&S_{n}^{(1)}\\[6.0pt] S_{n}^{(1)}&\frac{\gamma^{2n}-\gamma^{2}}{\gamma^{2}-1}\end{bmatrix}H^{(2)}_{n}$.
Then, taking the limit as $n\to\infty$, inequality (26) implies
$\varlimsup_{n\to\infty}\mathbb{E}[|X_{n}|^{2}] =lim$$=\varlimsup_{n\to\infty}\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]$ (45)
[…]
$=\frac{(\gamma^{2}-1)^{3}(L^{\prime}(\gamma^{-1}))^{2}}{\gamma^{4}L^{4}(\gamma^{-1})}+\frac{2(\gamma^{2}-1)^{2}}{\gamma}\frac{L^{\prime}(\gamma^{-1})}{L^{3}(\gamma^{-1})}+\frac{\gamma^{4}-1}{L^{2}(\gamma^{-1})},$
where the second equality follows from the definitions of $a_{n}$ and $b_{n}$ together with (43) and (44).
We next show that the right-hand side of (45), in the limit where $\gamma_{1}\to\gamma$ and $\gamma_{2}=\gamma\in\mathbb{R}$, coincides with the left-hand side of the power constraint (20). More precisely,
$\lim_{\begin{subarray}{c}\gamma_{1}\to\gamma\\ \gamma_{2}=\gamma\in\mathbb{R}\end{subarray}}\frac{1}{\Delta}\left(\frac{L_{11}}{\gamma_{2}^{2}}+\frac{L_{22}}{\gamma_{1}^{2}}+\frac{2L_{12}}{\gamma_{1}\gamma_{2}}\right)$ $=\frac{(\gamma^{2}-1)^{3}(L^{\prime}(\gamma^{-1}))^{2}}{\gamma^{4}L^{4}(\gamma^{-1})}+\frac{2(\gamma^{2}-1)^{2}}{\gamma}\frac{L^{\prime}(\gamma^{-1})}{L^{3}(\gamma^{-1})}+\frac{\gamma^{4}-1}{L^{2}(\gamma^{-1})}.$ (46)
This immediately establishes the power constraint (20) for Case B. The proof of the identity (46) is deferred to Appendix C. ∎
Now, combining Theorem 3.2 (the case $b=0$) and Theorem 4.1 (the case $b\neq 0$), we now present the main achievable-rate result of this paper for the SK(2) coding scheme, which provides a lower bound on $C_{\mathrm{SK2}(P)}$.
Corollary 4.2.
Suppose the noise $\{Z_{i}\}$ is an $AR(p)$ Gaussian process with parameters $\beta_{k}$, $|\beta_{k}|<1$ for all k=1,2,,pk=1,2,...,p, namely, it has the power spectral density
$S_{Z}(e^{i\theta})=|H_{Z}(e^{i\theta})|^{2}=\frac{1}{|L_{Z}(e^{i\theta})|^{2}}=\frac{1}{|\prod_{k=1}^{p}(1+\beta_{k}e^{i\theta})|^{2}}.$
Then, we have
$C_{\mathrm{SK2}}(P)\geq\max\left\{\bar{I}_{\mathrm{SK1}}(P),\bar{I}_{\mathrm{SK2}}(P)\right\},$
where $\bar{I}_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ are given by (12) and (19), respectively.
Proof.
This is an immediate consequence of the definition of $C_{\mathrm{FB}}(P)$, Theorem 3.2 (the case $b=0$), and Theorem 4.1 (the case $b\neq 0$) ∎
As the simplest application of Theorem 4.1 and Corollary 4.2, we consider the case of a flat noise spectrum, defined by
$S_{Z}(e^{i\theta})\equiv 1.$ (47)
This spectral density corresponds to a stationary noise process of the form $Z_{i}=W_{i},\penalty 10000\ i\geq 1$, where $\{W_{i}\}_{i=1}^{\infty}$ is a zero-mean, unit-variance white Gaussian process. This corresponds precisely to the AWGN channel (4). It follows directly from Theorem 3.2 and Corollary 4.2 that for the case $b=0$, the SK(2) scheme reduces to the SK(1) scheme and is therefore optimal for this channel, as noted in Remark 3.3. For the case $b\neq 0$, the following corollary demonstrates that, for such channel, the optimal pair $(\gamma_{1},\gamma_{2})$ in the optimization problem (19)-(20) satisfies $|\gamma_{1}|=|\gamma_{2}|$. Consequently, the SK(2) coding scheme yields a broader family of optimal coding scheme for the AWGN channel.
Corollary 4.3.
Consider the noise process $\{Z_{i}\}_{i=1}^{\infty}$ with the power spectral density given in (47). Then, the feedback capacity of the AWGN channel $Y_{i}=X_{i}+Z_{i},\penalty 10000\ i=1,2,$..., subject to the power constraint $P$, can be achieved via the SK(2) coding scheme defined in (15) and (14) for $b\neq 0$ with any parameter pair $(\gamma_{1},\gamma_{2})$ satisfying $|\gamma_{1}|=|\gamma_{2}|$.
Proof.
Note that for the AWGN channel (4) under consideration, we have $L_{Z}(e^{i\theta})\equiv 1$. Consequently, by applying Theorem 4.1, the optimization problem as formulated in (19), (20) and (21) simplifies such that the power constraint (20) reduces to
$\gamma^{2}_{1}\gamma^{2}_{2}-1\leq P,$ (48)
where $(\gamma_{1},\gamma_{2})$ denote the characteristic roots of equation (16) satisfying $|\gamma_{1}|>1$ and $|\gamma_{2}|>1$. The objective is therefore to maximize
$\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})\triangleq 2\log|\gamma_{1}|\wedge|\gamma_{2}|$
under the constraint (48). Next, we consider the two cases. First, suppose the characteristic roots $(\gamma_{1},\gamma_{2})$ satisfy $\gamma_{1}=\overline{\gamma_{2}}$ and $|\gamma_{2}|=\gamma>1$. In this case, $\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})=2\log|\gamma|$ and the power constraint (48) reduces to $\gamma^{4}-1\leq P$. Consequently,
$\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})\leq\frac{1}{2}\log(P+1)$
with equality if and only if $\gamma=(P+1)^{1/4}$, which corresponds to any pair $(\gamma_{1},\gamma_{2})$ such that $\gamma_{1}=\overline{\gamma_{2}}$ and $|\gamma_{2}|=(P+1)^{1/4}$.
For the second case, where $\gamma_{1},\penalty 10000\ \gamma_{2}\in\mathbb{R}$ with $\gamma_{1}\neq\gamma_{2}$, it follows from (48) that
$\max_{\begin{subarray}{c}\gamma_{1},\gamma_{2}\in\mathbb{R}\\ \gamma_{1}\neq\gamma_{2}\end{subarray}}\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})$ $=\max 2\log|\gamma_{1}|\wedge|\gamma_{2}|$
$=\frac{1}{2}\log(P+1),$
where the maximum is attained by any pair $(\gamma_{1},\gamma_{2})$ satisfying $\gamma_{1}=\pm(P+1)^{1/4}$ and $\gamma_{2}=\pm(P+1)^{1/4}$.
In summary, combining both cases we have shown that the maximum $\bar{I}_{\mathrm{SK2}}(P)$ in (19) is achieved whenever the characteristic roots $(\gamma_{1},\gamma_{2})$ satisfy $|\gamma_{1}|=|\gamma_{2}|=(P+1)^{1/4}$. In particular, the well-known feedback capacity of the AWGN channel, $\frac{1}{2}\log(P+1)$, is attained by the SK(2) coding scheme for the case $b\neq 0$ with any parameter pair $(\gamma_{1},\gamma_{2})$ satisfying $|\gamma_{1}|=|\gamma_{2}|$, as desired. ∎
We next consider the stationary AR(1) Gaussian channel $Y_{i}=X_{i}+Z_{i},\penalty 10000\ i\geq 1$, with noise spectrum
$S_{Z}(e^{i\theta})=\frac{1}{|1+\beta_{1}e^{i\theta}|^{2}}$
where $\beta_{1}\in(-1,1)$. This spectral density corresponds to the stationary noise process
$Z_{i}+\beta_{1}Z_{i-1}=W_{i},\penalty 10000\ \penalty 10000\ i\in\mathbb{Z}.$
According to [9] and [20], the feedback capacity of this channel under power constraint $P$ is known to be
$C_{\mathrm{FB}}(P)=-\log(x_{0}),$ (49)
where $x_{0}$ is the unique positive root of the fourth-order polynomial equation
$Px^{2}=\frac{1-x^{2}}{(1+|\beta_{1}|x)^{2}}.$ (50)
As noted in Remark 3.3, the SK(1) coding scheme defined in (8) and (9) is optimal for this channel. Specifically, we have
$C_{\mathrm{SK1}}(P)=C_{\mathrm{FB}}(P)=\bar{I}_{\mathrm{SK1}}(P),$
and thus by Corollary 4.2,
$C_{\mathrm{SK2}}(P)=C_{\mathrm{FB}}(P).$ (51)
This implies that the SK(2) coding scheme with $b=0$ is optimal for the stationary AR(1) Gaussian channel. The following example demonstrates, however, that for the case $b\neq 0$ the SK(2) coding scheme may not be optimal for AR(1) Gaussian channels unless $\beta_{1}=0$ (or equivalently, the AWGN channel (4)).
Example 4.4.
Based on Theorem 4.1 and the equations (49) and (50), we infer that
$\bar{I}_{\mathrm{SK2}}(P)\leq C_{\mathrm{FB}}(P),$
with equality if and only if $\beta_{1}=0$. For different values of $\beta_{1}$ and $P=1,5$, we plot $\bar{I}_{\mathrm{SK1}}(P)$, $\bar{I}_{\mathrm{SK2}}(P)$, and their difference as functions of $\beta_{1}$, as shown in Fig. 1.
Figure 1: Comparison of $\bar{I}_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ for stationary AR(1) Gaussian channels with $P=1,5$.
Thus, we have established the optimality of the SK(2) coding scheme for stationary AR(1) Gaussian channels, as summarized in the following corollary.
Corollary 4.5.
The SK(2) coding scheme achieves the feedback capacity of stationary AR(1) Gaussian channels.
Proof.
This follows directly from Corollary 4.2, Example 4.4 and (51). ∎
We now present the second example, which examines stationary AR(2) Gaussian channels and demonstrates that the SK(2) coding scheme can achieve strictly higher rates than the SK(1) coding scheme in certain scenarios.
Example 4.6.
Consider a stationary AR(2) Gaussian channel with noise power spectral density
$S_{Z}(e^{i\theta})=\frac{1}{|(1+\beta_{1}e^{i\theta})(1+\beta_{2}e^{i\theta})|^{2}},\penalty 10000\ \penalty 10000\ |\beta_{1}|,|\beta_{2}|<1.$
This spectrum corresponds to a stationary noise process governed by
$Z_{i}+(\beta_{1}+\beta_{2})Z_{i-1}+\beta_{1}\beta_{2}Z_{i-2}=W_{i},\penalty 10000\ \penalty 10000\ i\in\mathbb{Z}.$
To ensure real-valued AR coefficients, the pair $(\beta_{1},\beta_{2})$ must satisfy $\beta_{1}+\beta_{2}\in\mathbb{R}$ and $\beta_{1},\beta_{2}\in\mathbb{R}$. Under the power constraint $P$, Theorem 3.2 yields
$\bar{I}_{\mathrm{SK1}}(P)=-\log x_{\min},$
where $x_{\min}=\min_{x_{0}\in\mathcal{R}}|x_{0}|$ and $\mathcal{R}$ denotes the set of real roots of the polynomial equation
$Px^{2}=\frac{1-x^{2}}{((1+\beta_{1}x)(1+\beta_{2}x))^{2}}.$
Theorem 4.1 together with numerical results (Fig. 2) show that in certain cases
$\bar{I}_{\mathrm{SK2}}(P)>\bar{I}_{\mathrm{SK1}}(P).$
By Corollary 4.2, this implies
$C_{\mathrm{SK2}}(P)\geq\bar{I}_{\mathrm{SK2}}(P)>\bar{I}_{\mathrm{SK1}}(P).$
Thus, unlike for AR(1) Gaussian channels, the $\mathrm{SK}(1)$ coding scheme is not universally optimal. This disproves the conjecture of Butman [2] that in our notation $\bar{I}_{\mathrm{SK1}}(P)=C_{\mathrm{FB}}(P)$.
Figure 2: Comparison of $\bar{I}_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ for stationary AR(2) Gaussian channels with $\beta_{2}=0.5,\penalty 10000\ 0.9$ and $P=1$.
Appendix A Proof of Lemma 3.1
For any nonzero $x\in\mathbb{C}^{n}$, it holds that
$x^{\ast}(I_{n}+UU^{\ast})x$ $=x^{\ast}x+x^{\ast}UU^{\ast}x$
$=\|x\|^{2}+\|U^{\ast}x\|^{2}$
$\geq\|x\|^{2}$
$>0,$
which implies that $I_{n}+UU^{\ast}$ is positive definite and thus invertible. Then, we use the Woodbury matrix identity [6] to establish
$(I_{n}+UU^{\ast})^{-1}$ $=I_{n}-U(I_{m}+U^{\ast}U)^{-1}U^{\ast}$
$=I_{n}-UA^{-1}U^{\ast}$
and thus
$U^{\ast}(I_{n}+UU^{\ast})^{-1}U$ $=U^{\ast}U-U^{\ast}UA^{-1}U^{\ast}U$
$=A-I_{m}-(A-I_{m})A^{-1}(A-I_{m})$
$=I_{m}-A^{-1}.$
This further immediately implies that for $1\leq i,j\leq m$
$u_{i}^{\ast}(I_{n}+UU^{\ast})^{-1}u_{j}=\delta_{ij}-(A^{-1})_{ij},$ (52)
where $\delta_{ij}$ is the Kronecker delta funtion. Furthermore, since
$A^{-1}=\frac{1}{|A|}\text{adj}(A),$
where adj(A)(A) is the adjugate matrix of $A$, it holds that
$(A^{-1})_{ij}=(-1)^{i+j}\frac{1}{|A|}M_{ji},$
[…]
Appendix B Proof of Theorem 3.2
Consider the SK(1) coding scheme formulated by (8) and (9) over the stationary AR($p$) Gaussian channel (2) with the noise spectral density function $S_{Z}(e^{i\theta})$ given by (3). It then straightforward follows from (9) that
$V_{n}=\gamma^{n-1}U.$
Following the same arguments as used in (22), (23) and (24), we can express
$\widetilde{Y}_{n}=d_{n}U+W_{n}$
for $n\geq p+1$, where
$d_{n}=\gamma^{n-1}L(\gamma^{-1}).$ (53)
Thus, the mutual information between $V_{1}^{n}$ and $Y_{1}^{n}$ for $n\geq p+1$ is given by
$\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})$ $=\frac{1}{n}I(U;Y_{1}^{n})$ (54)
$=\frac{1}{n}I(U;\widetilde{Y}_{1}^{n})$
$\geq\frac{1}{n}I(U;\widetilde{Y}_{p+1}^{n})$
$=-\frac{1}{2n}\log\mathbb{E}\left[\left|U-\mathbb{E}[U|\widetilde{Y}_{p+1}^{n}\right|^{2}\right].$
Let
$D_{n}=\begin{bmatrix}d_{1}&d_{2}&...&d_{n}\end{bmatrix}^{T}.$
From (53) we obtain
$\mathbb{E}[|U-\mathbb{E}[U|\widetilde{Y}_{1}^{n}|^{2}]$ $=\frac{1}{1+\mathrm{Tr}(D_{n}D_{n}^{T})}$ (55)
$=\frac{1}{1+D_{n}^{T}D_{n}}$
$=\bigl[1+\frac{L^{2}(\gamma^{-1})}{\gamma^{2}-1}(\gamma^{2n}-1)\bigr]^{-1}.$
Substituting this into (54) and letting $n\to\infty$ yields (12), as desired.
Moreover, following a similar approach as in (26), we observe that for $n\geq p+2$
$\mathbb{E}[X_{n}^{2}]$ $=\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|Y_{1}^{n-1}]|^{2}]$ (56)
$\leq\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{p+1}^{n-1}]|^{2}]$
$=\gamma^{2(n-1)}\mathbb{E}[|U-\mathbb{E}[U|\widetilde{Y}_{p+1}^{n-1}]|^{2}],$
where the inequality follows because conditioning on a larger $\sigma$-algebra reduces the estimation error. Combining (55) with (56) yields
[…]
$=P.$
Finally, it follows from (3) that the power constraint (13) is satisfied, which completes the proof.
Appendix C Proof of (46)
For notational simplicity, let
$f(\gamma_{1},\gamma_{2})=\frac{1}{\Delta}\left(\frac{L_{11}}{\gamma_{2}^{2}}+\frac{L_{22}}{\gamma_{1}^{2}}+\frac{2L_{12}}{\gamma_{1}\gamma_{2}}\right).$
It then follows that
$f(\gamma_{1},\gamma_{2})=\frac{(\gamma_{1}^{2}-1)(\gamma_{2}^{2}-1)(\gamma_{1}\gamma_{2}-1)^{2}}{(\gamma_{1}-\gamma_{2})^{2}L^{2}(\gamma_{1}^{-1})L^{2}(\gamma_{2}^{-1})}g(\gamma_{1},\gamma_{2}),$ (57)
where
$g(\gamma_{1},\gamma_{2})=\frac{L^{2}(\gamma_{1}^{-1})}{\gamma_{1}^{2}-1}+\frac{L^{2}(\gamma_{2}^{-1})}{\gamma_{2}^{2}-1}-\frac{2L(\gamma_{1}^{-1})L(\gamma_{2}^{-1})}{\gamma_{1}\gamma_{2}-1}.$
Let $\gamma_{2}=\gamma\in\mathbb{R}$. The Taylor expansion of $g(\gamma_{1},\gamma)$ in $\gamma_{1}$ around $\gamma_{1}=\gamma$ gives:
$g($ $\gamma_{1},\gamma)=g(\gamma,\gamma)+\frac{\partial g}{\partial\gamma_{1}}(\gamma,\gamma)(\gamma_{1}-\gamma)+\frac{1}{2!}\frac{\partial^{2}g}{\partial\gamma_{1}^{2}}(\gamma,\gamma)(\gamma_{1}-\gamma)^{2}$ (58)
$+\frac{1}{3!}\frac{\partial^{3}g}{\partial\gamma_{1}^{3}}(\gamma,\gamma)(\gamma_{1}-\gamma)^{3}+.+\frac{1}{n!}\frac{\partial^{n}g}{\partial\gamma_{1}^{n}}(\gamma,\gamma)(\gamma_{1}-\gamma)^{n}+$.
Clearly,
$g(\gamma,\gamma)=0$ (59)
and
$\frac{\partial g}{\partial\gamma_{1}}(\gamma_{1},\gamma)=-\frac{2L(\gamma_{1}^{-1})L^{\prime}(\gamma_{1}^{-1})}{\gamma_{1}^{2}(\gamma_{1}^{2}-1)}-\frac{2\gamma_{1}L^{2}(\gamma_{1}^{-1})}{(\gamma_{1}^{2}-1)^{2}}+\frac{2L(\gamma^{-1})L^{\prime}(\gamma_{1}^{-1})}{\gamma_{1}^{2}(\gamma_{1}\gamma-1)}+\frac{2\gamma L(\gamma^{-1})L(\gamma_{1}^{-1})}{(\gamma_{1}\gamma-1)^{2}},$
which implies that
$\frac{\partial g}{\partial\gamma_{1}}(\gamma,\gamma)=0.$ (60)
Furthermore, we have
$\frac{\partial^{2}g}{\partial\gamma^{2}_{1}}(\gamma_{1},\gamma)$ $=\frac{2(L^{\prime}(\gamma_{1}^{-1}))^{2}}{\gamma_{1}^{4}(\gamma_{1}^{2}-1)}+\frac{2L(\gamma_{1}^{-1})L^{\prime\prime}(\gamma_{1}^{-1})}{\gamma_{1}^{4}(\gamma_{1}^{2}-1)}-\frac{2L^{2}(\gamma_{1}^{-1})}{(\gamma_{1}^{2}-1)^{2}}$
$+\frac{4(2\gamma_{1}^{2}-1)L(\gamma_{1}^{-1})L^{\prime}(\gamma_{1}^{-1})}{\gamma_{1}^{3}(\gamma_{1}^{2}-1)^{2}}+\frac{4L(\gamma_{1}^{-1})L^{\prime}(\gamma_{1}^{-1})}{\gamma_{1}(\gamma_{1}^{2}-1)^{2}}$
$-\frac{2L(\gamma^{-1})L^{\prime\prime}(\gamma_{1}^{-1})}{\gamma_{1}^{4}(\gamma_{1}\gamma-1)}-\frac{2(3\gamma_{1}\gamma-2)L(\gamma^{-1})L^{\prime}(\gamma_{1}^{-1})}{\gamma_{1}^{3}(\gamma_{1}\gamma-1)^{2}}$
$-\frac{2\gamma L(\gamma^{-1})L^{\prime}(\gamma_{1}^{-1})}{\gamma_{1}^{2}(\gamma_{1}\gamma-1)^{2}}-\frac{4\gamma^{2}L(\gamma^{-1})L(\gamma_{1}^{-1})}{(\gamma_{1}\gamma-1)^{3}}+\frac{8\gamma_{1}^{2}L^{2}(\gamma_{1}^{-1})}{(\gamma_{1}^{2}-1)^{3}},$
which implies that
$\frac{\partial^{2}g}{\partial\gamma_{1}^{2}}(\gamma,\gamma)=\frac{2\left(L^{\prime}(\gamma^{-1})\right)^{2}}{\gamma^{4}(\gamma^{2}-1)}+\frac{4L(\gamma^{-1})L^{\prime}(\gamma^{-1})}{\gamma(\gamma^{2}-1)^{2}}+\frac{2(\gamma^{2}+1)\left(L(\gamma^{-1})\right)^{2}}{(\gamma^{2}-1)^{3}}.$ (61)
Therefore, combining (57), (58), (59), (60) and (61) yields
$\lim_{\gamma_{1}\to\gamma}f(\gamma_{1},\gamma)$ $=\frac{(\gamma^{2}-1)^{4}}{2L^{4}(\gamma^{-1})}\frac{\partial^{2}g}{\partial\gamma_{1}^{2}}(\gamma,\gamma)$
$=\frac{(\gamma^{2}-1)^{3}}{\gamma^{4}}\frac{\left(L^{\prime}(\gamma^{-1})\right)^{2}}{L^{4}(\gamma^{-1})}+\frac{2(\gamma^{2}-1)^{2}}{\gamma}\frac{L^{\prime}(\gamma^{-1})}{L^{3}(\gamma^{-1})}+\frac{\gamma^{4}-1}{L^{2}(\gamma^{-1})},$
from which (46) follows immediately.